# A preconditioned difference of convex functions algorithm with extrapolation and line search

### v1

@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.26)
For the case 2 of (3.6), while the line search does not succeed with $N_{\max}$ searches, we assume $\lambda_{n}=0$ along with $\nu_{n}$ set to 0 and hence $x^{n+1}=\bar{x}^{n}$ as in (4.1), i.e.,
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\frac{1}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.27)
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)\leq f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-g_{2}\left(x^{n}\right)-\delta_{\lambda}^{n}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}$ (4.28)
$-\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$ (4.29)
For sufficiently large $N_{0}$, while $n>N_{0}$, we conclude $x^{n}\in B_{\varepsilon}(\bar{x}^{n-1})$ for small enough $\varepsilon$ by the square summability of $\{x^{n}\}_{n}$ as in (4.3) with $\bar{x}^{n-1}-x^{n-1}=(x^{n}-x^{n-1})/(1+\lambda_{n-1})$. Therefore, applying (4.21) as in the Assumption 1 with $x=x^{n}$, $\bar{x}=\bar{x}^{n-1}$, there exists $\bar{w}=\bar{w}^{n-1}$ and $\bar{w}^{n-1}\in\partial g_{1}\left(\bar{x}^{n-1}\right)$, such that
$g_{1}(x^{n})$ $\leq g_{1}(\bar{x}^{n-1})+\langle x^{n}-\bar{x}^{n-1},\bar{w}^{n-1}\rangle+\frac{\sigma}{2}\|x^{n}-\bar{x}^{n-1}\|^{2}$
$=\langle x^{n},\bar{w}^{n-1}\rangle-g_{1}^{*}(\bar{w}^{n-1})+\frac{\sigma}{2}\|x^{n}-\bar{x}^{n-1}\|^{2}.$ (4.30)
Substituting $g_{1}(x^{n})$ as in (4.30) into (4.29), we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)-\delta_{\lambda}^{n}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}$
$+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\delta_{\lambda}^{n}}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}-\frac{1-\delta_{\lambda}^{n}}{2}\left\|x^{n+1}-x^{n}\right\|_{M}^{2}.$
With $(x^{n}-x^{n-1})/(1+\lambda_{n-1})=\bar{x}^{n-1}-x^{n-1}$ and $\beta_{n}^{2}<\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}$, still with (4.6), reformulating the above inequality, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)$
$+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}-\delta_{n}^{\lambda}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}$
$+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{1}{2}\left(\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}-\beta_{n}^{2}\right)\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}.$
If taking $\eta$ be larger, whether $\lambda_{n}=0$ or $\lambda_{n}>0$, we have
$\delta_{n}^{\lambda}\left(\eta\cdot\lambda_{n}-\frac{\omega}{n+1}\right)\left\|\bar{x}^{n}-x^{n}\right\|^{2}+\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}\geq 0,$ (4.31)
since the left hand side of (4.31) is 0 for the case $\delta_{n}^{\lambda}=0$, i.e., $\lambda_{n}=0$ and $x^{n+1}=\bar{x}^{n}$. Then from Proposition 1, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{C_{\lambda}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}$
$\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}.$
Reformulating the above inequality, we obtain
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}+\frac{C_{\lambda}\mu_{\min}\left(M\right)}{2}\left\|x^{n}-x^{n-1}\right\|^{2}$
$\leq F\left(x^{n},\bar{w}^{n-1},x^{n-1}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n}-\bar{x}^{n-1}\right\|^{2}.$
Therefore, for $n>N_{0}$, we obtain
$H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\leq H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-C_{1}\left\|x^{n}-x^{n-1}\right\|^{2},$ (4.32)
@@
which means $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is decreasing monotonically. Furthermore, as we can see that $\bar{w}^{n}\in\partial g_{1}\left(\bar{x}^{n}\right)$, from Assumption 1, thus we obtain
$f\left(x^{n+1}\right)+\left<x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$=f\left(x^{n+1}\right)+g_{1}\left(\bar{x}^{n}\right)+\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$\geq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}+\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}-\frac{\sigma}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2}$
$\geq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right).$ (4.33)
We see that $f(x)+g(x)$ is bounded below, which implies
$\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}.$ (4.34)
@@
On the other hand, since $f(x)+g(x)$ is proper, closed, and level-bounded, we conclude that the iteration sequence $\{x^{n}\}_{n}$ is bounded with (4.33) and monotone decreasing of $H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)$.
@@
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)+\nabla g_{2}\left(x^{n_{i}}\right)\in\nabla f\left(\bar{x}^{n_{i}}\right)+\partial g_{1}\left(\bar{x}^{n_{i}}\right)-\nabla g_{2}(\bar{x}^{n_{i}}).$
@@
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla g_{2}\left(x^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)).$ (4.35)
@@
$\lim_{i\rightarrow\infty}\left\|\bar{x}^{n_{i}}-x^{n_{i}}\right\|\rightarrow 0\text{ and }\lim_{i\rightarrow\infty}\bar{x}^{n_{i}}=x^{*}.$ (4.36)
With $\bar{x}^{n_{i}}-x^{n_{i}}=(x^{n_{i}+1}-x^{n_{i}})/(1+\lambda_{n_{i}})$ and $\bar{x}^{n_{i}}-y^{n_{i}}=\bar{x}^{n_{i}}-x^{n_{i}}-\beta_{n_{i}}(x^{n_{i}}-x^{n_{i}-1})$, the Lipschitz continuity of $\nabla g_{2}$, and (4.36), we conclude that the formulation in (4.35) tends to zero while $n_{i}\rightarrow\infty$. We then obtain
$0\in\lim_{n_{i}\rightarrow\infty}-\partial g_{1}(\bar{x}^{n_{i}})+\nabla(f+g_{2})(\bar{x}^{n_{i}}).$
@@
$\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)=\left(\begin{array}[]{c}\nabla f\left(x^{n}\right)+\bar{w}^{n-1}-\nabla g_{2}\left(x^{n}\right)+\sigma\left(x^{n}-\bar{x}^{n-1}\right)\\ x^{n}-\partial g_{1}^{*}\left(\bar{w}^{n-1}\right)\\ -\sigma\left(x^{n}-\bar{x}^{n-1}\right)+M\left(\bar{x}^{n-1}-x^{n-1}\right)\\ -M\left(\bar{x}^{n-1}-x^{n-1}\right)\\ \end{array}\right).$ (4.37)
From first-order optimality condition, we have
$0\in\nabla f\left(\bar{x}^{n-1}\right)+\partial g_{1}\left(\bar{x}^{n-1}\right)-\nabla g\left(x^{n-1}\right)+M\left(\bar{x}^{n-1}-y^{n-1}\right).$
Since $\bar{\omega}^{n-1}\in\partial g_{1}\left(\bar{x}^{n-1}\right)$ and $\bar{x}^{n-1}\in\partial g_{1}^{*}\left(\bar{\omega}^{n-1}\right)$, with conditions $\bar{x}^{n-1}-x^{n-1}=\frac{x^{n}-x^{n-1}}{1+\lambda_{n-1}}$ and $x^{n}-\bar{x}^{n-1}=\lambda_{n-1}\left(\bar{x}^{n-1}-x^{n-1}\right)$, then reformulating the statement, we thus have
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\leq L\left\|x^{n}-\bar{x}^{n-1}\right\|+L_{g_{2}}\left\|x^{n}-x^{n-1}\right\|$
$+\left(2\sigma+1\right)\left\|x^{n}-\bar{x}^{n-1}\right\|+3B_{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|+B_{1}\left\|x^{n-1}-x^{n-2}\right\|\quad\left(*\right),$
where $B_{1}$ is the maximum singular value of $M$. Adjusting the expression above yields
$\left(*\right)\leq B_{2}\left\|x^{n}-x^{n-1}\right\|+B_{1}\left\|x^{n-1}-x^{n-2}\right\|\leq B\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right),$
where $B_{2}=(L+2\sigma+1)\cdot\lambda_{\max}+3B_{1}+L_{g_{2}}$ and $B=\max(B_{1},B_{2})$ are constants. Then we obtain the statement
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\leq B\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right).$ (4.38)
From (4.3), we know that $\lim_{n\rightarrow\infty}\left\|x^{n}-x^{n-1}\right\|\rightarrow 0$, we thus get
$\lim_{n\rightarrow\infty}\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)=0.$
Given the KL property of $H\left(x,w,\bar{y},y\right)$ with concave function $\psi\left(\cdot\right)$, and from (4.11), we establish
$\left[\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)-\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\zeta\right)\right]\cdot$
$\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)\geq\psi\prime\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)\cdot$
$\left[H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right]\cdot\mathrm{dist}\left(0,\partial H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)\right)$
$\geq H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\geq C_{1}\left\|x^{n}-x^{n-1}\right\|^{2}.$
Then we set
$\varPsi\left(x^{n-1},\bar{x}^{n-1},\bar{w}^{n-1},x^{n},\bar{x}^{n},\bar{w}^{n},x^{n+1},\bar{\zeta}\right)=\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-$
$\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right).$
@@
$\left\|x^{n}-x^{n-1}\right\|^{2}$ $\leq\tilde{B}\cdot\varPsi\left(x^{n-1},\bar{x}^{n-1},\bar{w}^{n-1},x^{n},\bar{x}^{n},\bar{w}^{n},x^{n+1},\bar{\zeta}\right)$
$\cdot\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right),$
@@
$\frac{1}{2}\left\|x^{n}-x^{n-1}\right\|$ $\leq\tilde{B}\cdot\left[\psi\left(H\left(x^{n},\bar{w}^{n-1},\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-\psi\left(H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right)\right]$
@@
$\frac{1}{2}\sum_{n=N_{0}}^{N}{\left\|x^{n}-x^{n-1}\right\|}\leq\frac{1}{4}\left(\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|-\left\|x^{N}-x^{N-1}\right\|\right)$
$+\tilde{B}\left[\psi\left(H\left(x^{N_{0}},\bar{w}^{N_{0}-1},\bar{x}^{N_{0}-1},x^{N_{0}-1}\right)-\bar{\zeta}\right)-\psi\left(H\left(x^{N+1},\bar{w}^{N},\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)\right].$
@@
$\lim_{N\rightarrow\infty}\psi\left(H\left(x^{N+1},\bar{w}^{N},\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)=\psi\left(0\right)=0.$
@@
$\sum_{n=N_{0}}^{\infty}\|x^{n}-x^{n-1}\|\leq$ $2\tilde{B}\cdot\psi(H\left(x^{N_{0}},\bar{w}^{N_{0}-1},\bar{x}^{N_{0}-1}x^{N_{0}-1}\right)-\bar{\zeta})+\frac{1}{2}\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|<\infty.$
Consequently, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$ and the global convergence of$\{x^{n}\}_{n}$. ∎
@@
Assume $\left\{x^{n}\right\}_{n}$ is generated from $pDCAe_{nls}$ in Algorithm 2 for solving (1.1). Define the perturbation energy function
$H\left(x,w,\bar{y},y\right)=f\left(x\right)+\left<x,w\right>-g_{1}^{*}\left(w\right)-g_{2}\left(x\right)+\frac{L\sigma}{2}\left\|x-\bar{y}\right\|^{2}+\frac{L}{2}\left\|\bar{y}-y\right\|_{M}^{2}.$
Suppose that $\left\{x^{n}\right\}_{n\in N}$ has a cluster $x^{*}$ and $H\left(x,w,\bar{y},y\right)$ is a KL function, then the following statements hold
@@
    Monotonicity: The sequence $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is monotonically decreasing and converges to some limit: $\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}$.

### v2

@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
Recalling the definition of $\widetilde{H}\left(\bar{x}^{n},x^{n}\right)$, then by $d^{n-1}=\bar{x}^{n-1}-x^{n-1}$, $\nu_{n-1}=\frac{\omega}{n}\left\|d^{n-1}\right\|^{2}$, and $\bar{x}^{n-1}-x^{n-1}=\frac{x^{n}-x^{n-1}}{1+\lambda_{n-1}}$, we have
$\widetilde{H}\left(\bar{x}^{n},x^{n}\right)=E\left(\bar{x}^{n}\right)+\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}\leq E\left(x^{n}\right)+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}$
$\leq E\left(\bar{x}^{n-1}\right)-\eta\lambda_{n-1}\left\|d^{n-1}\right\|^{2}+\nu_{n-1}+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}.$
Noticing that $\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)=E\left(\bar{x}^{n-1}\right)+\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}$, we thus obtain
$\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\geq\frac{1}{2}\left\|\bar{x}^{n-1}-x^{n-1}\right\|_{M}^{2}+\eta\lambda_{n-1}\left\|d^{n-1}\right\|^{2}-\nu_{n-1}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}$
$\geq\frac{1}{2\left(1+\lambda_{n-1}\right)^{2}}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}+\frac{\left(\eta\lambda_{n-1}-\frac{\omega}{n}\right)}{\left(1+\lambda_{n-1}\right)^{2}}\left\|x^{n}-x^{n-1}\right\|^{2}-\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}.$
Since $\eta\lambda_{n-1}-\frac{\omega}{n}>0$, and from the LSDE updates in Proposition 1, we have $\frac{1}{\left(1+\lambda_{n-1}\right)^{2}}-\beta_{n}^{2}>C_{\lambda}$. We thus obtain
$\widetilde{H}\left(\bar{x}^{n},x^{n}\right)-\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)\leq-\frac{C_{\lambda}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2},$ (4.23)
@@
which means $\left\{\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is decreasing monotonically. Furthermore, as we see that $f(x)+g(x)$ is bounded below, which implies
$\lim_{n\rightarrow\infty}\widetilde{H}\left(\bar{x}^{n},x^{n}\right)=\bar{\zeta}.$ (4.24)
@@
On the other hand, since $f(x)+g(x)$ is proper, closed, and level-bounded, we conclude that the iteration sequence $\{x^{n}\}_{n}$ is bounded with monotone decreasing of $\widetilde{H}\left(\bar{x}^{n},x^{n}\right)$.
@@
$M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)+\nabla g_{2}\left(x^{n_{i}}\right)\in\nabla f\left(\bar{x}^{n_{i}}\right)+\partial g_{1}\left(\bar{x}^{n_{i}}\right)-\nabla g_{2}(\bar{x}^{n_{i}}).$
@@
$M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla g_{2}\left(x^{n_{i}}\right)-\nabla g_{2}\left(\bar{x}^{n_{i}}\right)).$ (4.25)
@@
[…]
With $\bar{x}^{n_{i}}-x^{n_{i}}=(x^{n_{i}+1}-x^{n_{i}})/(1+\lambda_{n_{i}})$ and $\bar{x}^{n_{i}}-y^{n_{i}}=\bar{x}^{n_{i}}-x^{n_{i}}-\beta_{n_{i}}(x^{n_{i}}-x^{n_{i}-1})$, the Lipschitz continuity of $\nabla g_{2}$, and (4.26), we conclude that the formulation in (4.25) tends to zero while $n_{i}\rightarrow\infty$. We then obtain the following.
$0\in\lim_{n_{i}\rightarrow\infty}(-\partial g_{1}(\bar{x}^{n_{i}})+\nabla(f+g_{2})(\bar{x}^{n_{i}})).$
@@
$\partial\widetilde{H}\left(\bar{x}^{n},x^{n}\right)=\left(\begin{array}[]{c}\nabla f\left(\bar{x}^{n}\right)+\partial g_{1}\left(\bar{x}^{n}\right)-\nabla g_{2}\left(\bar{x}^{n}\right)+M\left(\bar{x}^{n}-x^{n}\right)\\ -M\left(\bar{x}^{n}-x^{n}\right)\\ \end{array}\right).$ (4.27)
From the first-order optimality condition, we have
$0\in\nabla f\left(\bar{x}^{n}\right)+\partial g_{1}\left(\bar{x}^{n}\right)-\nabla g_{2}\left(x^{n}\right)+M\left(\bar{x}^{n}-y^{n}\right).$ (4.28)
Substituting
$\partial g_{1}\left(\bar{x}^{n}\right)\in-\nabla f\left(\bar{x}^{n}\right)+\nabla g_{2}\left(x^{n}\right)-M\left(\bar{x}^{n}-y^{n}\right)$ (4.29)
with conditions $\bar{x}^{n-1}-x^{n-1}=\frac{x^{n}-x^{n-1}}{1+\lambda_{n-1}}$ and $x^{n}-\bar{x}^{n-1}=\lambda_{n-1}\left(\bar{x}^{n-1}-x^{n-1}\right)$, then reformulating the statement, we thus have
$\operatorname{dist}\left(0,\partial\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\right)\leq L_{g_{2}}\left\|\bar{x}^{n}-x^{n}\right\|+B_{1}\left\|\beta_{n}\left(x^{n}-x^{n-1}\right)\right\|+B_{1}\left\|\bar{x}^{n}-x^{n}\right\|$
$\leq\left(L_{g_{2}}+B_{1}\right)\left\|x^{n+1}-x^{n}\right\|+B_{1}\left\|x^{n}-x^{n-1}\right\|\leq B\left(\left\|x^{n+1}-x^{n}\right\|+\left\|x^{n}-x^{n-1}\right\|\right),$
where $B_{1}$ is the maximum singular value of $M$, $B_{2}=B_{1}+L_{g_{2}}$ and $B=\max(B_{1},B_{2})$ are constants. From (4.4), we see $\lim_{n\rightarrow\infty}\left\|x^{n}-x^{n-1}\right\|\rightarrow 0$. We thus arrive at
$\lim_{n\rightarrow\infty}\mathrm{dist}\left(0,\partial\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\right)=0.$
Given the KL property of $\widetilde{H}\left(\bar{x},x\right)$ with concave function $\psi\left(\cdot\right)$, and from (4.12), we establish
$\left[\psi\left(\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-\psi\left(\widetilde{H}\left(\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right)\right]\cdot\mathrm{dist}\left(0,\partial\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)\right)$
$\geq\psi\prime\left(\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)\cdot\left[\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\right]\cdot\mathrm{dist}\left(0,\partial\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)\right)$
$\geq\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\geq\frac{C_{\lambda}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}\geq C_{1}\left\|x^{n}-x^{n-1}\right\|^{2}.$
The existence of the positive constant $C_{1}$ is due to the positive definiteness of $M$. Then we set
$\varPsi\left(x^{n-1},\bar{x}^{n-1},x^{n},\bar{x}^{n},\bar{\zeta}\right)=\psi\left(\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-\psi\left(\widetilde{H}\left(\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right).$
@@
$\left\|x^{n}-x^{n-1}\right\|^{2}\leq\tilde{B}\cdot\varPsi\left(x^{n-1},\bar{x}^{n-1},x^{n},\bar{x}^{n},\bar{\zeta}\right)\cdot\left(\left\|x^{n}-x^{n-1}\right\|+\left\|x^{n-1}-x^{n-2}\right\|\right),$
@@
$\frac{1}{2}\left\|x^{n}-x^{n-1}\right\|\leq\tilde{B}\cdot\left[\psi\left(\widetilde{H}\left(\bar{x}^{n-1},x^{n-1}\right)-\bar{\zeta}\right)-\psi\left(\widetilde{H}\left(\bar{x}^{n},x^{n}\right)-\bar{\zeta}\right)\right]$
@@
$\frac{1}{2}\sum_{n=N_{0}}^{N}{\left\|x^{n}-x^{n-1}\right\|}\leq\tilde{B}\left[\psi\left(\widetilde{H}\left(\bar{x}^{N_{0}-1},x^{N_{0}-1}\right)-\bar{\zeta}\right)-\psi\left(\widetilde{H}\left(\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)\right]$
$+\frac{1}{4}\left(\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|-\left\|x^{N}-x^{N-1}\right\|\right).$
@@
$\lim_{N\rightarrow\infty}\psi\left(\widetilde{H}\left(\bar{x}^{N},x^{N}\right)-\bar{\zeta}\right)=\psi\left(0\right)=0.$
@@
$\sum_{n=N_{0}}^{\infty}{\parallel x^{n}}-x^{n-1}\parallel\leq 2\tilde{B}\cdot\psi\left(\widetilde{H}\left(\bar{x}^{N_{0}-1}x^{N_{0}-1}\right)-\bar{\zeta}\right)+\frac{1}{2}\left\|x^{N_{0}-1}-x^{N_{0}-2}\right\|<\infty.$
Consequently, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$ and the global convergence of $\{x^{n}\}_{n}$. ∎∎
@@
Assume the sequence $\left\{x^{n}\right\}_{n}$ is generated from $pDCAe_{nls}$ in Algorithm 2 for solving (1.1). Define the perturbation energy function
$H\left(\bar{x},x\right)=f\left(\bar{x}\right)+g_{1}\left(\bar{x}\right)-g_{2}\left(\bar{x}\right)+\frac{L}{2}\left\|\bar{x}-x\right\|_{M}^{2}.$
Suppose that $\left\{x^{n}\right\}_{n}$ has a cluster $x^{*}$ and $H\left(\bar{x},x\right)$ is a KL function, then the following statements hold
@@
    Monotonicity: The sequence $\left\{H\left(\bar{x}^{n},x^{n}\right)\right\}_{n>N_{0}}$ is monotonically decreasing and converges to some limit: $\lim_{n\rightarrow\infty}H\left(\bar{x}^{n},x^{n}\right)=\bar{\zeta}$.